OpenStudy (anonymous):

Solve the following problem. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit. The sum of three numbers is 26. The second number is twice the first and the third number is 6 more than the second. Find the numbers.

OpenStudy (texaschic101):

ok...I do not know what VEST is..... x is first number y is 2nd number z is 3rd number x + y + z = 26 2nd number is twice the first y = 2x 3rd number is 6 more then the second... z = 2x + 6 now it is just a matter of subbing... x + y + z = 26 x + (2x) + (2x + 6) = 26 -- combine like terms 5x + 6 = 26 5x = 26 - 6 5x = 20 x = 20/5 x = 4 y = 2x y = 2(4) y = 8 z = 2x + 6 z = 2(4) + 6 z = 8 + 6 z = 14 check.. x + y + z = 26 4 + 8 + 14 = 26 12 + 14 = 26 26 = 26 (correct) so your numbers are : 4,8,and 14